# FrameFusion: Combining Similarity and Importance for Video Token Reduction on Large Visual Language Models

### v1

@@ Abstract
[…]
@@ 1 Introduction
Large Vision-Language Models (LVLMs) have shown outstanding abilities across various video understanding scenarios, including temporal and spatial perception, recognition, and reasoning [34, 29, 14, 4]. Growing applications also demand LVLMs’ ability to process longer and more complex videos [21, 18, 24].
However, video understanding incorporates substantial computing overhead for LVLMs. Common LVLMs sample frames from the video, split each into image patches, and then sequentially embed them as vision tokens with a visual encoder. Though effective, it requires the processing of a huge amount of vision tokens for LVLMs. Take Google Gemini as an example, with the standard sampling rate of 1 frame per second (fps), it requires an astonishing amount of one million tokens to understand an hour’s video [21].
The importance-based token pruning has been extensively explored by previous works to solve the efficiency problem. They reduce vision tokens with various importance matrices, like accumulative attention score [23, 35, 7] and L2-norm of token feature [3]. However, there exist extensive important yet similar vision tokens in LVLMs due to frame resemblance and similar visual elements, while the effect of these tokens is less explored.
Figure 1: The central idea of FrameFusion. Compared with importance-based token pruning, FrameFusion additionally applies similarity-based token merging, keeping only important and unique vision tokens.
In this work, we recover similarity as a perpendicular property to importance for vision tokens reduction. The main idea is that although some frames are important for visual understanding, their high similarity leads to redundant information. These similar tokens, like those alike image patches from adjacent frames, hardly provide additional information, thus can be merged. To fully harness the potential of similarity-based merging, we study the token similarity distribution of LVLMs by answering three questions:
@@
    Where and when does similarity appear?
@@
    What is the value distribution of token similarity?
@@
    How consistent is the ranking of token similarity across layers?
Based on the observations, we propose FrameFusion to combine both similarity and importance for efficient LVLM token reduction. FrameFusion first evaluates the token similarity with minimum overhead and progressively merges similar tokens. After the similarity-based merging, FrameFusion then utilizes importance-based pruning to reduce tokens to the given computing budget.
Our contributions are three-folds:
  1. 1.
    We analyze the characteristics of token similarities across input positions and layers of LVLM.
  2. 2.
    We propose FrameFusion to hierarchically merge similar vision tokens and prune unimportant ones in a plug-and-play manner.
  3. 3.
    We conduct extensive experiments to show the effectiveness of FrameFusion on various video tasks.
Experiments show that FrameFusion pushes the Pareto front further for token compression. It successfully reduces the token budget and computing FLOPs to 30%, while maintaining comparable performance with the original dense model. The simple yet effective design helps FrameFusion to generalize across various model sizes, input lengths, and video tasks.
@@ 2.1 Large Vision Language Model (LVLMs)
The LVLM architecture typically consists of a vision encoder and a Large Language Model (LLM) [34, 29, 14, 4, 18, 24, 13]. The vision encoder converts visual inputs into token sequences, which the LLM then processes alongside text sequences to generate responses. Specifically, for video input, frames are first sampled temporally and then specially divided into sequences of image patches before sending to the vision encoder [34, 14, 29], as shown in Figure 1. Due to the high temporal and spatial resolution demands of complex video understanding tasks, token lengths can reach up to one million for an hour-long video [21], imposing significant computational overhead on LVLMs.
@@ 2.2 Token Compression
The first branch of work reduces the initial visual inputs before sending them to the vision encoder. They set rules to mix different temporal sampling frequencies [28, 34] and special resolutions [29, 5] when converting videos to input sequences, introducing trade-offs between visual detail and efficiency. Despite the simplicity, they inevitably incur direct detail losses and neglect the visual content guidance during compression.
Other works reduce tokens inside the vision encoder. They selectively retrieve [11] or condense [18] vision tokens in the vision encoder based on the guidance of text instructions. Yet, they require re-encoding all vision tokens if the text instruction changes, which incurs significant overheads for common multi-round conversation scenarios. Besides, an additional model fine-tuning is often needed to align the new vision encoding space.
Another branch of work focuses on token reduction in the subsequent LLM. For text-only tasks, previous works design static [8, 26, 10] or dynamic [12, 35, 9, 19] pruning pattern based on the importance of token (or KV-Cache). Emerging concurrent works highlight the specific token importance distribution for vision-language tasks [2, 16, 22, 33, 31], which further increase the sparsity of importance-based token pruning. However, as shown in Figure 4, token importance is inconsistent across different layers. It incurs prediction loss by pruning an unimportant token at shallow layers, which becomes important but inaccessible at deeper layers. FrameFusion falls in this category, exploring a more consistent and perpendicular token reduction method: similarity-based token merging.
@@ 3 Token Similarity Analysis
While token importance properties of LVLMs have been extensively studied [2, 16, 22, 33], a thorough and comparative analysis of token similarity remains lacking. Thus, we conduct oracle experiments to answer three key questions regarding the characteristics of token similarity, comparing them with token importance.
3.1 Experiment Setups and Notion Definitions
@@
In this section, we present the oracle results with the Llava-Video-7B [34] model tested on 128 video samples (64 frames at 1 fps) from the comprehensive VideoMME [6] datasets. All metrics are averaged over all test samples unless otherwise noted. Results on different models are similar and provided in Appendix 10.2.
@@
We define token importance and similarity using the notations $I^{(l)}\in\mathbb{R}^{N}$ and $S^{(l)}\in\mathbb{R}^{N}$, respectively, where $l$ indicate the layer index of the LLM. For clarity, we omit $l$ where it does not introduce ambiguity. We use the subscript $t$ to index the token along the input length dimension $N$. The token features at layer $l$ are represented as $\mathbf{X}^{l}\in\mathbb{R}^{N\times d}$.
@@
Following previous works [2, 35], the token importance $I^{(l)}_{t}$ is quantified using the cumulative attention score, calculated by summing the post-softmax attention scores vertically across the $t$-th column and averaging this sum across all attention heads at layer $l$.
@@
For token similarity $S^{(l)}_{t}$, we specifically define it as the cosine similarity between a vision token and its corresponding token from the preceding frame. The rationale for this definition is detailed in Section 3.2. Formally, if each frame is encoded into $P$ vision tokens, token similarity is calculated using the following equation:
$S_{t}=\frac{X_{t-P}^{T}X_{t}}{\lVert X_{t-P}\rVert_{2}\cdot\lVert X_{t}\rVert_{2}}$ (1)
3.2 Where and When Does Similarity Appear?
@@
Figure 2: Token similarities among all input tokens at the first LVLM layer in Llava-Video-7B models. For visual clarity, the color bar displays only the top 90% of similarity values. Vision tokens begin at index 14, with 210 tokens per frame.
We first investigate which tokens are most likely to demonstrate high similarity, as these are expected to have a greater potential for merging. To ensure figure clarity within the limited resolution, we limit the number of frames so that each input sequentially consists of 14 system prompt tokens, 840 vision tokens (representing 4 frames at 210 tokens per frame), and 20 user instruction tokens. Figure 2 displays the $N\times N$ cosine similarity matrix for all input tokens at the first LLM layer. A prominent 210th sub-diagonal appears, reflecting high similarities between tokens $i$ and $i+P$, where $P=210$ corresponds to the number of vision tokens per frame. This pattern underscores our initial observation:
@@
Observation 1. Spatially corresponding vision tokens between adjacent temporal frames exhibit higher cosine similarities than other token pairs.
Based on this observation, we focus on token similarities among these particular tokens and define token similarity as described in Equation 1. It is also worth noting that similar to token importance, the similarities between vision-language tokens are significantly lower than those between vision-vision or language-language tokens. This finding motivates a similar design choice to separate token compression for vision and language tokens, as proposed in prior work [2].
3.3 What Is the Distribution of Token Similarity?
@@
Figure 3: Heatmap of token similarity across different model layers. Each cell represents a similarity range at a specific layer, with color intensity denoting distribution frequency. The line overlay shows the average token similarity at each layer.
To decide the appropriate way for token compression, we explore the distribution of token similarities across different layers. As illustrated in Figure 3, although the average token similarity across layers remains relatively stable, there is a significant shift in the distribution of these similarities:
@@
Observation 2. High token similarities exhibit an obvious reduction in deeper model layers.
This trend is further confirmed by the variance in token similarity, which decreases from $6.0\times 10^{-2}$ at the first layer to $3.7\times 10^{-2}$ at the last layer. It illustrates a tightening of the distribution, with more numerical details in Appendix 10.1.
@@
The causal attention mechanism in LLMs contributes to this shift, where later vision tokens can aggregate information from earlier tokens, but not vice versa. The attention range differences lead to a divergence in the similarity of tokens that were initially similar, such as corresponding tokens from adjacent frames. This divergence accumulates as layers deepen.
@@
Based on this observation, FrameFusion prioritizes token merging at shallower layers rather than deeper layers to fully utilize the polarized high similarities.
3.4 How Consistent Is the Ranking of Token Similarity Across Layers?
The ranking consistency of token properties guides the compression strategy designs. For importance-based token pruning, the consistency across layers determines whether tokens deemed unimportant at shallow layers might become important at deeper layers. Early works [2, 23] adopt cascaded token pruning, where pruned tokens are no longer accessible in subsequent layers. These methods prioritize efficiency by eliminating token computations in both attention and Feed Forward Network (FFN) modules, and the importance is only computed once for pruned tokens. In contrast, more recent works [20, 17, 35] adopt non-cascaded pruning, which maintains the accessibility of all tokens across all layers by only pruning the KV-Cache at the attention layer. Yet, these methods do not reduce computation for the FFN and incur higher costs due to the need to compute importance for all tokens at every layer. Previous researches [8, 25] highlight the inconsistency of token importance across layers, which justifies re-evaluating pruning decisions at each layer. However, does a similar conclusion hold for token similarity?
@@ 3.4 How Consistent Is the Ranking of Token Similarity Across Layers?
Figure 4: Spearman Rank Correlation between adjacent layers for the Llava-Video-7B model.
To quantitatively assess the ranking consistency of token importance and token similarity, we calculate their respective Spearman Rank Correlations (SRC) [30] between adjacent layers. The SRC measures the degree to which the ranking of metrics in one layer correlates with the rankings in other layers. As shown in Figure 4, the SRCs for token similarity approach 1 across layers, indicating consistent ordering. In contrast, the SRCs for token importance are consistently lower, indicating weaker consistency across layers. We observe the following:
Observation 3. Token similarity demonstrates high rank consistency across different model layers.
Combining Observations 3 and 4, we conclude an interesting phenomenon: although the token similarity distribution condenses as layers deepen, tokens with the highest similarities remain the most similar. Motivated by this high consistency in similarity, FrameFusion employs similarity-based merging in a cascaded manner. Once highly-similar tokens are merged at shallow layers, they are not separated in subsequent computations.
@@
We further validate this cascaded design choice by quantifying the top-k retention rate across model layers. For notion simplicity, we define token uniqueness as 1 minus similarity. To calculate the top-k retention rate, we identify the set of top-$k$ tokens under token uniqueness or importance metrics at each layer. The top-k retention rate for layer $l$, starting from layer $i$, is defined as the intersection rate between the top-k tokens from both layers $l$ and $i$.
To formalize the calculation of the top-k retention rate, we denote the set of indices of the top-$k$ retaining tokens based on the scoring function $f:\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{N}$ at layer $l$ as follows:
@@
where scoring function $f$ can be token importance or token similarity, $\operatorname{Top-K}:\mathbb{R}^{N}\rightarrow\mathbb{N}^{K}$ function returns the indices of the top-k values. The top-k retention rate $R^{(l)}_{i}$ of layer $l$ with the starting layer $i$ is calculated as:
$R^{(l)}_{i}=|T^{(l)}\cap T^{(i)}|/|T^{(i)}|.$ (3)
As shown in Figure 5, we quantify the top-30% retention rate with shallow starting layers 00 to $2$ under different metrics. The token similarity exhibits a much higher retention rate than token importance. This finding validates the superiority of using token similarity over importance for cascaded token pruning, achieving a better accuracy trade-off.

### v2

# FrameFusion: Combining Similarity and Importance for Video Token Reduction on Large Vision Language Models

@@ Abstract
The increasing demand to process long and high-resolution videos significantly burdens Large Vision-Language Models (LVLMs) due to the enormous number of visual tokens. Existing token reduction methods primarily prune tokens based on importance metrics, such as cumulative attention scores. However, even important tokens may exhibit high redundancy caused by similarity among adjacent video frames and repetitive visual elements. To address this limitation, we propose FrameFusion, a novel token reduction approach integrating similarity-based merging with importance-based pruning. We conduct a thorough study on token similarity characteristics, revealing three key insights: (1) spatially corresponding visual tokens between adjacent frames have higher cosine similarities compared to other token pairs; (2) high token similarities prominently decrease in deeper model layers; and (3) token similarity rankings are highly consistent across different layers. Guided by these observations, FrameFusion computes token similarities exclusively between corresponding visual tokens from adjacent frames, applies token merging at initial successive layers followed by pruning in deeper layers, and adopts a cascaded merging strategy to further enhance efficiency. We evaluate FrameFusion comprehensively across six diverse LVLMs, ranging from 2B to 72B parameters, using five video benchmarks encompassing video retrieval, question-answering, and spatial-temporal understanding tasks. Experiments show that FrameFusion reduces visual tokens by 70%, achieving 1.6 – 3.6$\times$ end-to-end speedups, with an average performance impact of less than 3%. Our code is available at https://github.com/thu-nics/FrameFusion.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) have demonstrated remarkable capabilities across various video understanding tasks, including temporal and spatial perception, recognition, and reasoning [42, 34, 16, 5]. Increasingly demanding applications require LVLMs to process longer and more complex videos [27, 2, 21, 30].
However, handling extensive video data incurs substantial computational overhead. Typically, LVLMs sample frames from videos, divide each frame into image patches, and embed these sequentially as visual tokens through a visual encoder. While effective, this process generates an enormous number of tokens. For instance, Google’s Gemini, with a standard sampling rate of 1 frame per second (fps), needs to process approximately one million tokens to analyze an hour-long video [27].
Figure 1: The central idea of FrameFusion. Compared with importance-based token pruning, FrameFusion additionally applies similarity-based token merging, keeping only important and unique visual tokens.
Previous works primarily employ importance-based token pruning methods to mitigate efficiency demands. These approaches reduce visual tokens based on metrics such as cumulative attention scores [29, 43, 8] or normalized token feature [4]. Nevertheless, among the top 10% of tokens ranked by cumulative attention scores, 55% of them exhibit high redundancy, with a cosine similarity above 0.9 in the first layer of Llava-Video-7B.
In this work, we revisit token similarity as a perpendicular factor to token importance for reducing visual tokens. We argue that even important tokens can introduce redundancy due to visual similarities, particularly among adjacent frames. By merging these highly similar tokens, redundancy can be significantly reduced without compromising essential visual information.
To effectively utilize token similarity, we first systematically investigate its characteristics within LVLMs. We find that (1) token similarity predominantly occurs between spatially corresponding tokens from adjacent frames, (2) similarities exhibit high values particularly at shallow layers, and (3) the token similarity rankings are highly consistent across layers.
Based on these insight, we propose FrameFusion, a plug-and-play approach that integrates similarity-based merging with importance-based pruning. FrameFusion efficiently computes token similarities, progressively merges similar tokens at shallow layers, and subsequently applies importance-based pruning to adhere to computational constraints. Our contributions are summarized as follows:
@@
    We systematically analyze token similarity characteristics across input positions and layers in LVLMs.
@@
    We propose FrameFusion, a novel, post-training method that integrates similarity-based token merging with importance-based pruning for video token reduction.
@@
    We validate the effectiveness of FrameFusion through extensive experiments across diverse LVLMs, model sizes, input lengths, and video benchmarks.
Experiments confirm that FrameFusion effectively advances the Pareto front in token compression. It reduces visual tokens to 30%, achieving 1.6–3.6$\times$ end-to-end speedup while maintaining less than a 3% performance drop over dense models. Its simple and effective design ensures broad applicability across various tasks and scenarios.
@@ 2.1 Large Vision Language Model (LVLMs)
The LVLM architecture typically consists of a visual encoder and a Large Language Model (LLM) [42, 34, 16, 5, 21, 30, 15]. The visual encoder converts visual inputs into token sequences, which the LLM then processes alongside text sequences to generate responses. Specifically, for video input, frames are first sampled temporally and then specially divided into sequences of image patches before sending to the visual encoder [42, 16, 34], as shown in Figure 1. Due to the high temporal and spatial resolution demands of complex video understanding tasks, token lengths can reach up to one million for an hour-long video [27], imposing significant computational overhead on LVLMs.
@@ 2.2 Token Compression
The first branch of work reduces the initial input before sending them to visual encoder. They set rules to mix different temporal sampling frequencies [33, 42] and special resolutions [34, 6] when converting videos to input sequences, introducing trade-offs between visual detail and efficiency. Despite simplicity, they inevitably incur direct detail losses and neglect the content guidance for compression.
Other works reduce tokens inside the visual encoder. They selectively retrieve [12] or condense [21] visual tokens in the visual encoder. Yet, they require re-encoding all visual tokens if the text instruction changes, which incurs significant overheads for common multi-round conversation scenarios. Besides, an additional model fine-tuning is often needed to align the new visual encoding space [13].
Another branch of work focuses on token reduction in the subsequent LLM. For text-only tasks, previous works design static [9, 31, 11] or dynamic [14, 43, 10, 22] pruning pattern based on the importance of token (or KV-Cache). Emerging concurrent works highlight the specific token importance distribution for vision-language tasks [3, 18, 28, 41, 37], which further increase the sparsity of importance-based token pruning. However, as shown in Figure 4, token importance is inconsistent across different layers. It incurs prediction loss by pruning an unimportant token at shallow layers, which becomes important but inaccessible at deeper layers. FrameFusion falls in this category, exploring a more consistent and perpendicular token reduction method: similarity-based token merging. More related works are discussed in Appendix 11.
@@ 3 Token Similarity Analysis
While token importance in LVLMs has been extensively explored [3, 18, 28, 41, 39], the characteristics of token similarity remain under-investigated. To bridge this gap, we conduct comprehensive oracle experiments analyzing token similarity and contrasting it with token importance.
3.1 Experimental Setup and Definitions
@@
In this section, we present oracle experiment results on the Llava-Video-7B [42] model using the first 128 video samples from the VideoMME dataset [7], each comprising 64 frames sampled at 1 fps. Metrics reported are averaged over all samples unless otherwise noted. Similar results on other models are included in Appendix 10.
@@
We define token importance and token similarity as $I^{(l)}\in\mathbb{R}^{N}$ and $S^{(l)}\in\mathbb{R}^{N}$, respectively, with $l$ indicating the LLM layer index and $N$ denoting the number of input tokens. We use subscript $t$ to index the token along the input length dimension $N$. For simplicity, we omit the layer index when contextually clear. Input hidden features for layer $l$ are denoted as $\mathbf{X}^{(l)}\in\mathbb{R}^{N\times d}$.
@@
Following previous works [3, 43], token importance $I^{(l)}_{t}$ is computed using the cumulative attention score, calculated by summing the post-softmax attention scores vertically across the $t$-th column and averaging this sum across all attention heads at layer $l$.
@@
For token similarity $S_{t}^{(l)}$, we define it specifically between each visual token and its spatially corresponding token from the preceding frame, based on our empirical observation detailed in Section 3.2. Formally, token similarity is computed with cosine similarity:
$S_{t}=\frac{X_{t-P}^{T}X_{t}}{\lVert X_{t-P}\rVert_{2}\cdot\lVert X_{t}\rVert_{2}},$ (1)
where $P$ represents the number of visual tokens per frame, and $T$ represents the matrix transpose.
3.2 Where Does High Similarity Occur?
@@
Figure 2: Token similarities among all input tokens at the first LVLM layer in Llava-Video-7B models. For visual clarity, the color bar displays only the top 90% of similarity values. Visual tokens begin at index 14, with 210 tokens per frame.
We first analyze which tokens typically exhibit high similarity, as these tokens have greater potential for merging. To maintain visual clarity, we limit the video frames to 4 (resulting in 840 visual tokens at $P=210$ tokens per frame), with 14 system prompt tokens and 20 user instruction tokens before and after the visual tokens. Figure 2 illustrates the $N\times N$ cosine similarity matrix at the first LVLM layer. A distinct 210th sub-diagonal emerges, highlighting significantly higher similarities between tokens at positions $i$ and $i+P$. Statistically, the average cosine similarity of 0.62 between these tokens far exceeds the average similarity (0.28) observed elsewhere. Hence, we conclude:
@@
Observation 1. Spatially corresponding visual tokens from adjacent frames exhibit higher cosine similarity compared to other token pairs.
Based on this observation, we focus on similarities among these particular tokens and define token similarity as described in Equation 1. Additionally, sub-diagonals at multiples of 210 tokens also show relatively high similarities. This indicates that visual redundancy extends across multiple consecutive frames, which can also be identified by sequentially examining adjacent frames.
3.3 What Is the Token Similarity Distribution Across Layers?
@@
Figure 3: Heatmap of token similarity across model layers. Each cell represents a similarity range at a specific layer, with color intensity denoting distribution frequency. The line overlay shows the mean token similarity per layer.
To identify the most effective layers for token compression, we analyze the distribution of token similarity values across model layers. Figure 3 shows that, although the mean token similarity remains relatively stable across layers, the distribution shifts noticeably:
@@
Observation 2. High token similarity values decrease in deeper model layers.
This trend is particularly evident in Llava-Video model. While the mean similarity decreases only marginally (from 0.62 to 0.60), the distribution significantly condenses as layers deepen. Specifically, the 30th percentile similarity value decreases from 0.90 at the first layer to 0.72 at the last. Additional numerical results are in Appendix 10.1.
@@
The causal attention mechanism in LLMs contributes to this shift because tokens at later positions can aggregate information from earlier tokens, but not vice versa. This directional aggregation causes initially similar tokens, such as those corresponding across adjacent frames, to diverge increasingly at deeper layers.
@@
Given this observation, FrameFusion prioritizes token merging at shallower layers to effectively leverage these initially high similarities.
3.4 Is Token Similarity Ranking Consistent Across Layers?
The consistency of token similarity rankings across layers determines whether tokens that are similar at shallow layers remain similar in deeper layers.
@@ 3.4 Is Token Similarity Ranking Consistent Across Layers?
Figure 4: Spearman Rank Correlation (SRC) between adjacent layers for the Llava-Video-7B model.
We first quantify the consistency of token similarity rankings between adjacent layers using Spearman Rank Correlation (SRC) [35], which measures the correlation of rankings across different layers. For comparison, we also compute the SRC for token importance. As shown in Figure4, token similarity maintains consistently high SRC values approaching 1, indicating stable rankings across layers. In contrast, token importance shows lower and unstable SRC values, suggesting that tokens considered unimportant at shallow layers might become important at deeper layers.
@@
We further examine ranking consistency between shallow and deep layers using the top-k retention rate. For notion simplicity, we define token uniqueness as 1 minus similarity. To calculate the retention rate, we identify the top-$k$ tokens at each layer based on token uniqueness or importance. The top-k retention rate at layer $l$ with respect to a starting layer $i$ is defined as the intersection ratio between the top-k tokens at layers $l$ and $i$. Formally:
@@
where $f$ represents token uniqueness or token similarity, and $\operatorname{Top-K}$ selects indices of the top-k values. The retention rate $R^{(l)}i$ is then calculated as:
$R^{(l)}{i}=|T^{(l)}\cap T^{(i)}|/|T^{(i)}|.$ (3)
As shown in Figure 5, we calculate the top-30% retention rate at shallow starting layers 00 to $2$ under different metrics. The token similarity exhibits a much higher retention rate than token importance. Based on these analyses, we conclude:
Observation 3. Token similarity rankings are more consistent across layers than token importance rankings.
Combining Observations 3 and 5, we highlight a key phenomenon: while the similarity between highly similar tokens decreases at deeper layers, the relative similarity rankings of these tokens remain stable. Motivated by this stable ranking, FrameFusion merges highly similar tokens at shallow layers, maintaining this reduction throughout subsequent layers.